OpenStudy (anonymous):

Evaluate the line integral f(x,y)ds along the parametric curve: f(x,y)=3x+4y^2;x=2t-1, y=t+5, -1(

OpenStudy (amistre64):

\[ds =\sqrt{(x')^2+(y')^2}\]

OpenStudy (amistre64):

well, in this case i spose fx and fy for partials

OpenStudy (anonymous):

so i take the partial of both x and y with respect to t and plus that under a square root next to the function f(x,y) and integrate?

OpenStudy (amistre64):

yes, but dont forget to replace x and y with their t equations within f f(x,y)=3x+4y^2;x=2t-1, y=t+5 f(t)=3(2t-1)+4(t+5)^2

OpenStudy (amistre64):

or am i mixing things up?

OpenStudy (anonymous):

ds = |v(t)| dt

OpenStudy (amistre64):

im mixing a vector feild in with this i believe

OpenStudy (anonymous):

no i believe you are right, let me write it out real quick

OpenStudy (anonymous):

you're fine @amistre64

OpenStudy (anonymous):

so i would just get a sqrt(5) from the ds, then the function would read integral of (3(2t-1)+4(t+5)^2)dt (times the sqrt(5))?

OpenStudy (amistre64):

..... thnx ;)

OpenStudy (amistre64):

yep, but put the sqrt5 inside the integral

OpenStudy (anonymous):

yes correct, @razroz

OpenStudy (anonymous):

thank you so much

OpenStudy (amistre64):

good luck ;)